Solve the system by substitution. 3x – y = –7
x + y = –9

Answer by stanbon(48545)

Solve the system by substitution. 1st: 3x – y = –7
2nd: x + y = –9
-----------
1st: y=3x+7
Substitute in 2nd to get:
x+3x+7=-9
4x=-16
x=-4
Substutute into 2nd to solve for y:
-4+y=-9
y=-5
Check in 1st as follows:
3(-4)-(-5)=-7
-12+5=-7
-7=-7
Cheers,
Stan H.